. Find the general solution to = y-t+(y-t) +1 and draw a graph of the solutions in the ty-plane
Answer #1

(4-7)+(y-)+1; DE Lets suppose ly-t= x 0 Put o 2 in D.E. dol +) = x +x+) dx = x+0² dx = at x(x1 Integrate it both sides 166 )* = * 1-ety 1+ etty You = elite), = मक [ = y-t. Recally e D 9-t- 13- हमा | ye) = t + रस final egn.Ay Y=tt altt y=flt) at (110) (0-1)